# Predictability of Performance in Communication Networks Under Markovian Dynamics

### v3

@@ V-A Approximation and Upper Bound Validation
To assess the effectiveness of our proposed approximation method in Proposition 1 and the upper bound in Theorem 2, we analyze a Geo/Geo/1/K queue with default parameters $\mu=0.5$, $\rho=0.85$, and $K=128$. Figure 4 illustrates the posterior and marginal delay distributions given various observed queue lengths for this queue.
@@
In Figure 5, we validate the approximation method by comparing it to the exact predictability derivation presented in Theorem 1. The main finding is that the accuracy of the approximation is significantly influenced by the disparity between the observed state’s forecast distribution and the marginal distribution. As illustrated in Figure 4, states that are closer to the marginal distribution display less accurate approximations in Figure 5, whereas states further from the marginal distribution (those above 16) demonstrate more accurate approximations.
@@ V-B Geo/Geo/1/K Dynamics and Predictability
To investigate the predictability in systems with Geo/Geo/1/K queues further, next, we examine the impact of the observed state and utilization factor ($\rho$). In the following analysis, we obtain predictability using Theorem 3 and Proposition 1.
@@
Figure 8: Predictability of a single Geo/Geo/1 queue, under different configurations. Solid lines illustrate numerical derivation and dashed lines the approximation in Proposition 1.
@@
Figure 8 demonstrates the predictability when $\mu=0.4$, $K=100$, and the observed state $x$ is either 3, 9, or 15 times greater than the expected state denoted by $\chi$. It is evident from the figure that the sojourn time is more predictable when there is a significant backlog or when the queue length is substantially different from the expected queue length. This implies that the queue remains predictable for a certain duration into the future, and we have sufficient information to predict the delay distribution for a period of time before it converges to the stationary delay distribution. In real systems, especially those operating in high utilization regimes where efficiency is critical, this analysis highlights the importance of monitoring the queue. High utilization often corresponds to more significant backlog situations, where predictability is maintained for longer durations.
@@
V-C Predictability of Multi-Hop Geo/Geo/1 Queues
In this part, we evaluate 2 multi-hop scenarios under different levels of observability. In the first case, illustrated in Figure 10, we evaluate predictability for a system that consists of 5 tandem Geo/Geo/1/K queues, all with the same service rate $\mu=0.4$ and highly congested with the observed state of $15\chi$. We consider 3 different configurations: observations that include all 5 queues, observations including only the first 3 queues, and observations including only the first queue. We used Equations 37 and 38to derive exact predictability curves shown via solid lines. As was expected, predictability decreases as our ability to observe the system declines. Then the dashed line curves show the upper bounds derived using Lemma 5 and Theorem 3 (UB). In the first case where all queues are observable, we derived the bound by summing up the predictability of all hops. However, when it comes to the case with only one observable queue, the bound equals the predictability of that queue. This is because if a component is not observable, its predictability is zero and will not contribute to the sum in Equation 39. Therefore, we have $D_{n}(L)\leq D^{(1)}_{n}(L)$, which constitutes the upper bound for the partially observable system.
@@ V-C Predictability of Multi-Hop Geo/Geo/1 Queues
Figure 10: Predictability of a 5-hop tandem Geo/Geo/1 queuing system under different observability levels. All queues have the same service rate of $\mu=0.4$ and $\lambda=0.32$
@@
The next scenario we analyze in multi-hop networks involves a queue that becomes a bottleneck and how its observability affects predictability. Since this queue will contribute more to the end-to-end delay, if it is not observable, we expect the predictability to decrease significantly. In a setup, we examine 3 tandem Geo/Geo/1/K queues with an arrival rate of $\lambda=0.34$, where the middle queue has a lower service rate, i.e., the service rates are $\mu^{(1)}=0.4$, $\mu^{(2)}=0.38$, and $\mu^{(3)}=0.4$ respectively. Suppose $\chi^{(m)}$ as the expected queue length of hop $m$. All three queues are again highly congested with the observed states of 10 times their expected queue size. As shown in Figure 11, we consider 4 different configurations: the observations include all 3 queues, the observations include only the first queue, the observations include only the last queue, and the observation include only the bottleneck queue. Note that in tandem M/M/1 queues, when the system is in equilibrium, regardless of the service rates, all queues will have the same arrival rate as the first queue [37]. Using that, Equation 37, and Equation 38, we derive exact predictability curves shown via solid lines. The figure indicates that the observation of one of the less utilized queues provides similar benefits in terms of predictability. However, observing these queues alone results in significantly lower predictability compared to the scenario in which only the bottleneck queue is observable. The predictability of the latter case is very close to that of complete observability. Similar to Figure 10, the dashed lines represent the upper bound derived using Lemma 5 and Theorem 3 (UB) and it is less tight when the predictability of all hops is summed up. Note that $\chi^{(1)}=\chi^{(3)}$ due to similar arrival and service rates.
@@
Figure 11: Predictability of a 3-hop tandem Geo/Geo/1/K queuing system with a bottleneck queue in the middle. Service rates are $0.4$, $0.38$, and $0.4$ respectively. Arrival rate: $0.34$.
@@ VI Conclusions
In this paper, we presented a definition of performance predictability for systems with Markovian conditions, with partial and imperfect observability. We applied this framework to Geo/Geo/1 queuing models in both single-hop and multi-hop contexts. We illustrated the practical implications of studying predictability in these systems. Our evaluations highlight the effectiveness of the proposed definitions in modeling crucial trade-offs in future communication systems incorporating QoS predictors. For example, we illustrated the overhead-predictability tradeoff and demonstrated how a lack of observability can deteriorate predictability.
@@
The behavior of the underlying Markov processes greatly impacts the predictability of the system. Drawing on the literature on Markov chain mixing times, including spectral gap analysis, provides insights into how random system conditions transition. We examined the connection between system predictability and Markov chain behavior by employing an upper limit on predictability derived from spectral analysis methods. Future research can investigate conductance analysis to further improve our comprehension of Markov chain dynamics. A possible direction for future research is to identify or detect hidden Markov models in the environment with which the communication network interacts. Incorporating these models into predictability analysis will offer insights into the feasibility and timescales of QoS predictions.
In future communication systems where performance prediction is essential, it is important to improve predictability efficiently. This involves enhancing observability and directing monitoring efforts towards critical points within the network, such as performance bottleneck nodes or queues. Additionally, it is necessary to identify whether the environment or application’s behavior is predictable and subsequently learn it for performance prediction. By incorporating these insights, more efficient communication systems can be achieved.
@@ Appendix A Proof of Theorem 2
In this proof, our approach is to bound total variation distance using chi-squared distance. For arbitrary probability distributions $p$ and $q$, according to Pinsker’s inequality, total variation distance is upper bounded as
@@ Appendix B Proof of Proposition 1
We start from Theorem 3, Equation 42 where in the inner sum over $k$, the term $\alpha^{L}_{k}$ converges to zero, while other terms remain nonzero, but only when $L$ is large, and except when $k/(K+1)$ is small. Therefore, we simplify the rest of the terms for the case when $k/(K+1)$ is small. As for the $\sin$ terms, as $k/(K+1)$ is small, we approximate
$\sin\left(\frac{(x+1)k\pi}{K+1}\right)\approx\sin\left(\frac{xk\pi}{K+1}\right),\\ \sin\left(\frac{(y+1)k\pi}{K+1}\right)\approx\sin\left(\frac{yk\pi}{K+1}\right)$ (B.1)
@@
$D_{n}(L)=\frac{1}{2}\sum_{z}\biggl{|}\sum_{y=1}^{K}\frac{2}{K+1}\beta^{\frac{y-x}{2}}(1-\sqrt{\beta})^{2}\\ \sum_{k=1}^{K}\frac{\left(\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\cos(\frac{k\pi}{K+1})\right)^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \sin\left(\frac{xk\pi}{K+1}\right)\sin\left(\frac{yk\pi}{K+1}\right)\text{NB}(z;y,\mu)\biggr{|}.$ (B.2)
@@
$\sum_{y=1}^{K}\beta^{y/2}\sin{\left(\frac{yk\pi}{K+1}\right)}\text{NB}(z;y,\mu)\approx\\ \sqrt{\beta}\frac{\sin{\left(\frac{k\pi}{K+1}\right)}}{1-2\sqrt{\beta}\cos{\left(\frac{k\pi}{K+1}\right)}+\beta}\text{NB}(z;1,\mu)$ (B.3)
@@
Next we simplyfy $\alpha_{k}^{L}$ term using the binomial approximation for large $L$:
$(a+b\cos(c))^{L}\approx(a+b)^{L}\left(1-\frac{bc^{2}}{2(a+b)}\right)^{L}\approx\\ (a+b)^{L}e^{-L\frac{bc^{2}}{2(a+b)}}$
Substitute $\alpha_{k}^{L}$ with above approximation will give us:
$\left(\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\cos(\frac{k\pi}{K+1})\right)^{L}\approx\\ \left(\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}\right)^{L}e^{-L\kappa(k\pi/(K+1))^{2}}$ (B.4)
@@
$\kappa=\frac{\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}{\lambda\mu+\bar{\lambda}\bar{\mu}+2\sqrt{\lambda\mu\bar{\lambda}\bar{\mu}}}$

### v4

@@ V-B Approximation and Upper Bound Validation
Next, we shift our focus to the queuing system and analyze its predictability. To assess the effectiveness of our proposed approximation method in Proposition 1 and the upper bound in Theorem 2, we analyze a Geo/Geo/1/K queue with default parameters $\mu=0.5$, $\rho=0.85$, and $K=128$. Figure 6 illustrates the posterior and marginal delay distributions given various observed queue lengths for this queue.
@@
In Figure 7, we validate the approximation method by comparing it to the exact predictability derivation presented in Theorem 1. From the exact calculations we observe that predictability decreases as the lead time $L$ increases. Regarding the approximation, the main finding here is that its accuracy is significantly influenced by the disparity between the observed state’s forecast distribution and the marginal distribution. As illustrated in Figure 6, states that are closer to the marginal distribution display less accurate approximations in Figure 7, whereas states further from the marginal distribution (those above 16) demonstrate more accurate approximations.
@@
This widening gap highlights a limitation of Theorem 2 in this specific scenario. The theorem provides a general upper bound based on the spectral properties of the transition matrix. However, this bound, while mathematically valid, is derived under worst-case assumptions regarding convergence to the stationary distribution. For larger queue sizes, the sojourn time predictability is primarily determined by transitions within the set of higher queue length states. A large queue tends to stay large for a while, due to the traffic intensity ($\rho=0.85$). The general spectral bound, which considers all possible transitions, becomes less representative of these specific, dominant dynamics. It overestimates the influence of transitions that are less relevant when starting from a large queue, leading to a looser bound. In contrast, for smaller queue sizes, a wider range of transitions are relevant to the short-term and medium-term behavior, and the spectral bound provides a tighter approximation.
@@ V-C Geo/Geo/1/K Dynamics and Predictability
To investigate the predictability in systems with Geo/Geo/1/K queues further, next, we examine the impact of the observed state and utilization factor ($\rho$). In the following analysis, we obtain predictability using Theorem 3 and Proposition 1. Furthermore, to benchmark our approach against a standard information-theoretic measure, we also quantify predictability using cross-entropy. We define the cross-entropy predictability by:
$D^{\text{CE}}_{n}(L)=-\sum_{z}\Pr(Z_{n+L})\log\left(\frac{\Pr(Z_{n+L}\mid O_{0:n}=o_{0:n})}{\Pr(Z_{n+L})}\right).$ (53)
(a) TV distance predictability
(b) Cross entropy (CE) predictability
@@
Figure 10: Predictability of a single Geo/Geo/1/K queue, under different configurations. Solid lines illustrate exact numerical derivation and dashed lines the approximation in Proposition 1.
@@
Figure 10 illustrates the predictability of the Geo/Geo/1/K system, quantified using two distinct measures: TV distance (Sub-figure 10(a)) and cross-entropy predictability (Sub-figure 10(b)). The results are presented for parameters $\mu=0.4$, $K=100$ and the observed state $x$ is either 3, 9, or 15 times greater than the expected state denoted by $\chi$. Focusing on the TV-based predictability, it is evident from the figure that the sojourn time is more predictable when there is a significant backlog or when the queue length is substantially different from the expected queue length. This implies that the queue remains predictable for a certain duration into the future, and we have sufficient information to predict the delay distribution for a period of time before it converges to the stationary delay distribution. In real systems, especially those operating in high utilization regimes where efficiency is critical, this analysis highlights the importance of monitoring the queue. High utilization often corresponds to more significant backlog situations, where predictability is maintained for longer durations.
Comparing the TV distance (Figure 10(a)) and cross-entropy predictability (Figure 10(b)), reveals key differences arising from their mathematical definitions. The TV distance, being a true metric, is bounded between 0 and 1. This provides a standardized scale for assessing the degree of predictability, clearly indicating whether, and to what extent, the system’s future state is influenced by its initial condition. Cross-entropy, conversely, is unbounded and can reach arbitrarily large values as observed in Figure 10(b). This sensitivity of cross-entropy to discrepancies in low-probability regions contrasts with the more moderate response of the bounded TV distance. Furthermore, the established connections between TV distance and Markov chain mixing times offer advantages in terms of analytical tractability, as demonstrated by the derivation of Theorem 2 in this work.
Figure 11 shows the predictability of delay in a Geo/Geo/1/K queueing system as a function of different maximum queue sizes (K), with fixed arrival and service parameters ($x=3\chi$, $\rho=0.85$). The figure shows that larger queue sizes (higher $K$) result in significantly higher predictability. This demonstrates that systems with more buffering capacity retain information about future delays for longer, making them more predictable. When $K$ is small, the queue has a restricted capacity for holding customers. As the offered load increases while $K$ remains limited, new arrivals are more frequently blocked or lost. This leads to a lower average queue size, as the system cannot accumulate long queues and the system reaches its maximum capacity quickly, and past information becomes less relevant. The curves for larger $K$ values show higher predictability since the current queue size provides more information about the future system state. A longer queue now is a stronger indicator of longer delays in the future, and that information persists for a longer lead time. However, we observe a decreasing marginal gain as $K$ increases.
Figure 11: Predictability of a single Geo/Geo/1/K queue under different maximum queue sizes, derived using exact numerical calculations.
Building on the analysis of maximum queue size $K$, we now focus on Proposition 2, where performance is packet loss due to blockage rather than sojourn time. Figure 12 shows numerical results for a Geo/Geo/1/$K$ queue under various buffer sizes and observed queue lengths. When predicting near-future packet drops, an observation that the queue is nearly full (dashed lines) leads to higher predictability compared to an empty queue (solid lines). Further, when the queue is empty at present, having a smaller buffer increases predictability at short lead times: fewer packets are needed to overflow the system, so either “no blocking” or “rapid filling” scenarios become more pronounced and predictable. Conversely, if the queue is full and $K$ is large, there is a strong likelihood that blocking persists in the short run, making it comparatively easier to forecast a block. As the lead time $L$ grows, however, this ordering can invert. A larger $K$ may ultimately offer more predictable long-term behavior after the system evolves closer to its steady state: with more capacity, the Markov chain dynamics can settle into patterns where blocking occurs less erratically.
Figure 12: Predictability of packet loss in a single Geo/Geo/1/K queue under different maximum queue sizes and observed queue lengths, derived using exact numerical calculations in Proposition 2.
@@
V-D Predictability of Multi-Hop Geo/Geo/1 Queues
In this subsection, we first analyze a multi-hop scenario focusing solely on the impact of the number of hops on system predictability, followed by two multi-hop scenarios with varying levels of observability.
In the first case, shown in Figure 14, we examine tandem Geo/Geo/1/K queues with an increasing number of hops, while keeping the arrival rate ($\alpha=0.32$) and service rate ($\mu=0.4$) the same for all queues. Counterintuitively, the predictability increases as the number of hops increases. This seemingly paradoxical result stems from the fact that we are observing all queues in the system ($x=[15\chi,15\chi,...]$) and the target metric is the total sojourn time through the entire system. With a single hop, the sojourn time is highly variable. However, as the number of hops increases, and each hop’s state is known, the total sojourn time becomes the sum of multiple, independent (due to the memoryless property of the Geo/Geo/1 queue) sojourn times. By the law of large numbers, the sum of these independent random variables tends towards a more predictable, less variable value. Essentially, the individual fluctuations in each queue’s sojourn time tend to average out over the entire multi-hop path, leading to a more stable, and therefore more predictable, total sojourn time. While this reasoning offers a plausible initial explanation, future work should formally investigate and validate this proposed mechanism.
@@ V-D Predictability of Multi-Hop Geo/Geo/1 Queues
Figure 14: Predictability of multi-hop tandem Geo/Geo/1/K queuing systems under different number of hops. All queues have the same service rate of $\mu$ = 0.4 and $\alpha$ = 0.32
As illustrated in Figure 15, we evaluate predictability for a system that consists of 5 tandem Geo/Geo/1/K queues, all with the same service rate $\mu=0.4$ and highly congested with the observed state of $15\chi$. We consider 3 different configurations: observations that include all 5 queues, observations including only the first 3 queues, and observations including only the first queue. We used Equations 38 and 39 to derive exact predictability curves shown via solid lines. As was expected, predictability decreases as our ability to observe the system declines. Then the dashed line curves show the upper bounds derived using Lemma 5 and Theorem 3 (UB). In the first case where all queues are observable, we derived the bound by summing up the predictability of all hops. However, when it comes to the case with only one observable queue, the bound equals the predictability of that queue. This is because if a component is not observable, its predictability is zero and will not contribute to the sum in Equation 40. Therefore, we have $D_{n}(L)\leq D^{(1)}_{n}(L)$, which constitutes the upper bound for the partially observable system.
Figure 15: Predictability of a 5-hop tandem Geo/Geo/1 queuing system under different observability levels. All queues have the same service rate of $\mu=0.4$ and $\alpha=0.32$
@@
The next scenario we analyze in multi-hop networks involves a queue that becomes a bottleneck and how its observability affects predictability. Since this queue will contribute more to the end-to-end delay, if it is not observable, we expect the predictability to decrease significantly. In a setup, we examine 3 tandem Geo/Geo/1/K queues with an arrival rate of $\alpha=0.34$, where the middle queue has a lower service rate, i.e., the service rates are $\mu^{(1)}=0.4$, $\mu^{(2)}=0.38$, and $\mu^{(3)}=0.4$ respectively. Suppose $\chi^{(m)}$ as the expected queue length of hop $m$. All three queues are again highly congested with the observed states of 10 times their expected queue size. As shown in Figure 16, we consider 4 different configurations: the observations include all 3 queues, the observations include only the first queue, the observations include only the last queue, and the observation include only the bottleneck queue. Note that in tandem M/M/1 queues, when the system is in equilibrium, regardless of the service rates, all queues will have the same arrival rate as the first queue [55]. Using that, Equation 38, and Equation 39, we derive exact predictability curves shown via solid lines. The figure indicates that the observation of one of the less utilized queues provides similar benefits in terms of predictability. However, observing these queues alone results in significantly lower predictability compared to the scenario in which only the bottleneck queue is observable. The predictability of the latter case is very close to that of complete observability. Similar to Figure 15, the dashed lines represent the upper bound derived using Lemma 5 and Theorem 3 (UB) and it is less tight when the predictability of all hops is summed up. Note that $\chi^{(1)}=\chi^{(3)}$ due to similar arrival and service rates.
@@
Figure 16: Predictability of a 3-hop tandem Geo/Geo/1/K queuing system with a bottleneck queue in the middle and 3 imperfect observation scenarios. Service rates are $0.4$, $0.38$, and $0.4$ respectively. Arrival rate: $0.34$.
@@ VI Conclusions
In this paper, we presented a definition of performance predictability for systems with Markovian conditions, with partial and imperfect observability. We applied this framework to Geo/Geo/1 queuing models in both single-hop and multi-hop contexts, as well as to a random-walk-based model representing the connection state of the vehicles traversing in a V2X environment. We illustrated the practical implications of studying predictability in these systems. Our evaluations highlight the effectiveness of the proposed definitions in modeling crucial trade-offs in future communication systems, including V2X applications, incorporating QoS predictors. For example, we illustrated the overhead-predictability tradeoff and demonstrated how a lack of observability can deteriorate predictability.
@@
The behavior of the underlying Markov processes greatly impacts the predictability of the system. Drawing on the literature on Markov chain mixing times, including spectral gap analysis, provides insights into how random system conditions transition. We examined the connection between system predictability and Markov chain behavior by employing an upper limit on predictability derived from spectral analysis methods. Future research can explore conductance analysis to enhance our understanding of Markov chain dynamics. A possible direction for future research is to identify or detect hidden Markov models in the environment with which the communication network interacts. Incorporating these models into predictability analysis will offer insights into the feasibility and timescales of QoS predictions.
In future communication systems where performance prediction is essential, it is important to improve predictability efficiently. This involves enhancing observability and directing monitoring efforts towards critical points within the network, such as performance bottleneck nodes or queues. Additionally, it is necessary to identify whether the environment or application’s behavior is predictable and subsequently learn it for performance prediction. Incorporating these insights can lead to more efficient communication systems.
@@ Appendix A Proof of Theorem 2
In this proof, our approach is to bound TV distance using chi-squared distance. For arbitrary probability distributions $p$ and $q$, according to Pinsker’s inequality, total variation distance is upper bounded as
@@ Appendix B Proof of Proposition 1
We start from Theorem 3, Equation 43 where in the inner sum over $k$, the term $\gamma^{L}_{k}$ converges to zero, while other terms remain nonzero, but only when $L$ is large, and except when $k/(K+1)$ is small. Therefore, we simplify the rest of the terms for the case when $k/(K+1)$ is small. As for the $\sin$ terms, as $k/(K+1)$ is small, we approximate
$\sin\left(\frac{(x+1)k\pi}{K+1}\right)\approx\sin\left(\frac{xk\pi}{K+1}\right),\\ \sin\left(\frac{(y+1)k\pi}{K+1}\right)\approx\sin\left(\frac{yk\pi}{K+1}\right).$ (B.1)
@@
$D_{n}(L)=\frac{1}{2}\sum_{z}\biggl{|}\sum_{y=1}^{K}\frac{2}{K+1}\beta^{\frac{y-x}{2}}(1-\sqrt{\beta})^{2}\\ \sum_{k=1}^{K}\frac{\left(\alpha\mu+\bar{\alpha}\bar{\mu}+2\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}\cos(\frac{k\pi}{K+1})\right)^{L}}{\left[1-2\sqrt{\beta}\cos\left(\frac{k\pi}{K+1}\right)+\beta\right]}\\ \sin\left(\frac{xk\pi}{K+1}\right)\sin\left(\frac{yk\pi}{K+1}\right)\text{NB}(z;y,\mu)\biggr{|}.$ (B.2)
@@
$\sum_{y=1}^{K}\beta^{y/2}\sin{\left(\frac{yk\pi}{K+1}\right)}\text{NB}(z;y,\mu)\approx\\ \sqrt{\beta}\frac{\sin{\left(\frac{k\pi}{K+1}\right)}}{1-2\sqrt{\beta}\cos{\left(\frac{k\pi}{K+1}\right)}+\beta}\text{NB}(z;1,\mu).$ (B.3)
@@
Next we simplyfy $\gamma_{k}^{L}$ term using the binomial approximation for large $L$:
$(a+b\cos(c))^{L}\approx(a+b)^{L}\left(1-\frac{bc^{2}}{2(a+b)}\right)^{L}\approx\\ (a+b)^{L}e^{-L\frac{bc^{2}}{2(a+b)}}.$
Substitute $\gamma_{k}^{L}$ with above approximation will give us:
$\left(\alpha\mu+\bar{\alpha}\bar{\mu}+2\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}\cos(\frac{k\pi}{K+1})\right)^{L}\approx\\ \left(\alpha\mu+\bar{\alpha}\bar{\mu}+2\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}\right)^{L}e^{-L\kappa(k\pi/(K+1))^{2}},$ (B.4)
@@
$\kappa=\frac{\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}}{\alpha\mu+\bar{\alpha}\bar{\mu}+2\sqrt{\alpha\mu\bar{\alpha}\bar{\mu}}}.$
@@
Appendix C Random Walk Parameter Extraction
We analyzed a real-world dataset comprising LTE measurements collected from multiple CSV files introduced in [48], each containing time-series entries for attributes such as NetworkMode, CQI, Speed, and DL_bitrate. For each file, we retained rows that corresponded to operator A, used the LTE mode, and featured valid CQI readings. Downlink bitrates (originally in kbps) were rounded to the nearest discrete value in Mbps. After discarding missing or insufficient data, we combined the cleaned results across all files for further analysis.
(a) CQI = $3$
(b) CQI = $12$
Figure 17: Empirical vs. fitted negative binomial distributions for two CQI values and measured downlink LTE throughput in Mbps from dataset introduced in [48].
Next, we fit a Negative Binomial (NB) distribution to the downlink throughput for each CQI level $\{1,\dots,15\}$ as we illustrate two examples in Figure 17. Whenever a CQI level had sufficiently many samples (e.g., more than 10 valid observations), we used the method of moments to estimate the NB parameters $\bigl{(}r,p\bigr{)}$. We then performed a linear regression across the CQI levels to relate the estimated parameters $\{\,r,\,p\}$ to CQI in the form
$r(\mathrm{CQI})\;=\;\alpha_{r}\,\mathrm{CQI}+\beta_{r},\quad p(\mathrm{CQI})\;=\;\alpha_{p}\,\mathrm{CQI}+\beta_{p},$
and saved both the raw estimates and the regression results for subsequent simulation. The regressions results are depicted in Figure 18.
(a) Negative binomial parameter $p$ estimation via $\alpha_{p}\,\mathrm{CQI}+\beta_{p}$
(b) Negative binomial parameter $r$ estimation via $\alpha_{r}\,\mathrm{CQI}+\beta_{r}$
Figure 18: Linear regression results to determine the relationship between the parameters of the negative binomial distribution and different CQI values.
To capture how CQI evolves over time, we employed a lazy random walk model on a finite state space $\{1,\dots,15\}$, in which the CQI remains unchanged between consecutive time steps with probability $p$, and transitions to adjacent states with probability $\frac{1-p}{2}$. We estimated $p$ empirically by counting how often $X_{n+1}=X_{n}$ in the dataset. To distinguish between vehicular (mobile) and near-static connections, we applied a speed threshold (e.g., 20 km/h) and computed separate values of $p$. In the dataset at hand, vehicular connections yielded $p\approx 0.6$, whereas near-static connections yielded $p\approx 0.9$. These results reflect the differing degrees of CQI fluctuation in mobile versus almost-static environments. Thus, with the estimated parameters, we construct the conditional distribution and define the corresponding Markov chain, enabling the application of the predictability framework.